OpenStudy (anonymous):

If i took 3 times y how do i get the answer?

OpenStudy (saifoo.khan):

\[3 \times y = 3y\]

OpenStudy (anonymous):

how would i gwet that explain?

OpenStudy (anonymous):

pretend y is ANY number. say 6 for example. then 3x6 = 18. it's basic multiplication

OpenStudy (inkyvoyd):

y+y+y

OpenStudy (inkyvoyd):

=2*y+y

OpenStudy (inkyvoyd):

=3*y

OpenStudy (inkyvoyd):

3y=3*y, if that's what's confusing you.

OpenStudy (inkyvoyd):

If you need an example, I'd be happy to show you.

OpenStudy (anonymous):

What is y??????

OpenStudy (anonymous):

y is a variable, you can substitute any number in for a variable, in this case y is that variable.